OpenStudy (anonymous):

x + y = 3k x - y = k The solution of the system shown is _____. (0, 2k) (2k, 0) (2k, k)

OpenStudy (anonymous):

first of all you have to use substitution method x + y = 3k x - y = k ------------------------------------------- x cancels out and it becomes 2y = 2k so the y intercept will be just k so the answer is the last one

OpenStudy (anonymous):

If you use the addition method you can cancle out the y's first and have 2x=4k

OpenStudy (anonymous):

then divide 2 from both sides to get x=2k

OpenStudy (anonymous):

finally chose either problem to subsitute 2k for x to find why. I'll use the fist one. so 2k+y=3k

OpenStudy (anonymous):

subtract 2k from both sides to get y=k In this case your answer would be (2k,k)